### v1

# Detecting LLM Hallucination Through Layer-wise Information Deficiency: Analysis of Unanswerable Questions and Ambiguous Prompts

Abstract
Large language models (LLMs) frequently generate confident yet inaccurate responses, introducing significant risks for deployment in safety-critical domains. We present a novel approach to detecting model hallucination through systematic analysis of information flow across model layers when processing inputs with insufficient or ambiguous context. Our investigation reveals that hallucination manifests as usable information deficiencies in inter-layer transmissions. While existing approaches primarily focus on final-layer output analysis, we demonstrate that tracking cross-layer information dynamics ($\mathcal{L}$I) provides robust indicators of model reliability, accounting for both information gain and loss during computation. $\mathcal{L}$I improves model reliability by immediately integrating with universal LLMs without additional training or architectural modifications.
Detecting LLM Hallucination Through Layer-wise Information Deficiency: Analysis of Unanswerable Questions and Ambiguous Prompts
Hazel Kim, Adel Bibi, Philip Torr, Yarin Gal University of Oxford {hazel.kim, yarin.gal}@cs.ox.ac.uk, {adel.bibi, philip.torr}@eng.ox.ac.uk
$\dagger$\$$\dagger$\$footnotetext: Correspondence to Hazel Kim <hazel.kim@cs.ox.ac.uk>.$\ddagger$\$$\ddagger$\$footnotetext: Equal advising.
1 Introduction
Large language models (LLMs) have achieved unprecedented success across diverse natural language tasks, particularly in complex reasoning ranging from commonsense to arithmetic knowledge (Achiam et al., 2023; Touvron et al., 2023; Abdin et al., 2024). However, these models face a critical challenge known as hallucination—a phenomenon where responses appear convincingly authoritative despite being inaccurate (Ji et al., 2023; Xu et al., 2024b; Liu et al., 2023). While numerous empirical studies have investigated potential sources of hallucination, recent theoretical work by Xu et al. (2024a) demonstrates the fundamental impossibility of eliminating this issue through any computable function.
Following Xu et al. (2024a), hallucination can be formally defined as the failure of LLMs to accurately reproduce the desired output of a computable function. This theoretical framework establishes that hallucination is an inherent characteristic of LLMs, persisting regardless of architectural choices, learning algorithms, prompting strategies, or training data composition. Building on this theoretical foundation, we hypothesize that hallucination emerges when LLMs lack sufficient information necessary for their computational functions to transmit messages across their internal processing systems effectively.
Figure 1: Our proposed $\mathcal{L}$I correlates more strongly with question difficulty as influenced by answerability and prompt clarity—compared to final layer entropy (i.e., $\mathcal{V}$-usable information). The AUROC shows model classification in (un)answerable questions. ’With prompts’ are binary-type instruction prompts appended to the question (see Table 1).
(a) $\mathcal{V}$-usable information ($\mathcal{V}$I) per each layer. The information is not monotonically increased or decreased as the layer goes deeper. The original estimation of $\mathcal{V}$-usable information applies to the final layer (i.e., final layer $\mathcal{V}$I in Fig. 1) as illustrated in red.
(b) Cumulative $\mathcal{V}$I from the beginning to the layer $i$. The final layer tends to increase the overall scores, and sometimes exceeds those of others. It is crucial to cumulate the quantities up to the final layer to represent $\mathcal{L}$I, shown as the last quantities on x-axis. Relevant cases are further illustrated in Fig. 6.
Figure 2: $\mathcal{L}$I Estimation. The reason we consider the overall layer information rather than just the final layer is that it provides a more comprehensive understanding of the model’s behavior. The final layer $\mathcal{V}$I shows almost contradictory results from $\mathcal{L}$I; the CoQA has the lowest final layer $\mathcal{V}$I and the highest $\mathcal{L}$I, for example.
Prior research has focused primarily on analyzing final outputs to assess LLM confidence (Osband et al., 2023; Ahdritz et al., 2024; Lin et al., 2024) or identifying inherent data ambiguities that lead to predictive uncertainty (Cole et al., 2023; Kuhn et al., 2023a). However, the fundamental internal mechanisms of LLMs remain under-explored. Our analysis reveals that uncertainty estimation based solely on output layers or final computations overlooks critical insights into model self-confidence, thereby limiting our ability to detect hallucinatory behaviors.
To investigate LLM internal mechanisms, we build upon recent information theory frameworks proposed by Xu et al. (2020) and Ethayarajh et al. (2022). Their work introduces the concept of $\mathcal{V}$-usable information—the quantity of information a model family $\mathcal{V}$ can utilize to predict $Y$ given $X$. This metric indicates prediction difficulty: lower $\mathcal{V}$-usable information corresponds to more challenging predictions for $\mathcal{V}$.
While these findings are significant, we empirically demonstrate that $\mathcal{V}$-usable information provides sub-optimal insight into model self-confidence because it only applies to the final layer. We propose layer-wise usable information ($\mathcal{L}$I), which quantifies the information changes within certain layers and aggregates those information dynamics across all model layers. As shown in Fig. 1, $\mathcal{L}$I provides more reliable indicators of LLM performance than final layer $\mathcal{V}$-usable information ($\mathcal{V}$I), particularly in detecting subtle variations in instruction prompt effectiveness.
Prior work on $\mathcal{V}$-usable information established that computation can create usable information during feature extraction, constituting a violation of the data processing inequality (DPI) in information theory (LeCun et al., 2015; Xu et al., 2020). Our research extends this understanding by demonstrating that LLMs both create and lose usable information during layer-wise updates. As illustrated in Fig.2(a), information flow is non-monotonic across layers, highlighting the limitation of analyzing only the final layer’s computation. The proposed $\mathcal{L}$I framework captures this dual nature of DPI compliance and violation inherent to LLMs, with Fig.2 revealing substantial divergence between $\mathcal{L}$I and $\mathcal{V}$-usable information patterns. An additional explanation of DPI is presented in Section 3.2.
Prior work by Hewitt et al. (2021) has extended probing techniques on layers. However, their investigation remains limited in scope, primarily comparing layer-wise information against baseline performance and exploring part-of-speech information within the models.
To evaluate our framework, we focus on scenarios where LLMs must respond to queries with constrained contextual information (i.e., unanswerable questions), a setting particularly prone to hallucination. Our case study in Table 1 demonstrates that $\mathcal{L}$I strongly correlates with question difficulty—as influenced by answerability and prompt clarity—compared to $\mathcal{V}$I, which shows no significant correlation with model predictions.
Contributions
Our primary contributions are:
  • We propose $\mathcal{L}$I as a superior detector of unanswerable questions compared to existing baselines (Section 5.2), without requiring architectural modifications or additional training.
  • We demonstrate that $\mathcal{L}$I effectively captures model confidence across varying levels of task difficulty induced by different instruction prompts (Section 5.3).
  • We interpret that comprehensive layer tracking provides better insights into model internal confidence than single-layer analysis, using either initial or final layer (Sections 3 and 5.4).
Context
Once there was a beautiful fish named Asta. Asta lived in the ocean. There were lots of other fish in the ocean where Asta lived. They played all day long. One day, a bottle floated by over the heads of Asta and his friends. They looked up and saw the bottle. "What is it?" said Asta’s friend Sharkie. "It looks like a bird’s belly," said Asta. But when they swam closer, it was not a bird’s belly. It was hard and clear, and there was something inside it. The bottle floated above them. They wanted to open it. They wanted to see what was inside. So they caught the bottle and carried it down to the bottom of the ocean. They cracked it open on a rock. When they got it open, they found what was inside. It was a note. The note was written in orange crayon on white paper. Asta could not read the note. Sharkie could not read the note. They took the note to Asta’s papa. "What does it say?" they asked. Asta’s papa read the note. He told Asta and Sharkie, "This note is from a little girl. She wants to be your friend. If you want to be her friend, we can write a note to her. But you have to find another bottle so we can send it to her." And that is what they did.
Question Prompt Label Ans. $\mathcal{V}$I $\mathcal{L}$I Difficulty
What was the name of the fish? Binary question "Is this answerable?" Yes. $\checkmark$ $0.3$ $0.2$
Open-ended question w/ "don’t know" option Asta. $\times$ $0.3$ $-0.7$**
No prompts Asta. $\checkmark$ 1.2 $-6.8$****
Were they excited? Binary question "Is this answerable?" No. $\checkmark$ $0.3$ $-0.4$*
Open-ended question w/ "don’t know" option Don’t know. $\times$ $-0.3$ $-1.4$***
No prompts Unknown. $\times$ $0.2$ $-7.7$*****
Table 1: An example from CoQA. The $\mathcal{L}$I scores show more comprehensive overview of difficulties in questions and prompts than $\mathcal{V}$I. Ans.: Answers generated by a language model. If the generated answer is correct, then $\checkmark$. Otherwise, $\times$. $\mathcal{V}$I: $\mathcal{V}$-usable information only applied to the final layer. $\mathcal{L}$I: Layer-wise usable information accumulated from the beginning to the final layers (our proposed method). Difficulty: the level of difficulty estimated by average performance scores. It also matches with human intuition that expects to be more difficult in questions without instruction prompts than those with binary prompts, for example.
2 Background
In this section, we explain the information-theoretic foundations for measuring model-usable information, established by Xu et al. (2020) and Ethayarajh et al. (2022). Consider a model family $\mathcal{V}$ that maps text input $X$ to output $Y$. The $\mathcal{V}$-usable information quantifies the amount of information a model family can extract to predict $Y$ given $X$. This metric inversely correlates with prediction difficulty. The lower the $\mathcal{V}$-usable information, the harder the dataset is for $\mathcal{V}$. Consider text that is encrypted or translated into a language with more complex grammatical structures. Such transformations decrease $\mathcal{V}$-usable information, making the prediction of $Y$ given $X$ more challenging within the same model family $\mathcal{V}$.
This concept challenges traditional information theory principles, notably Shannon’s mutual information (Shannon, 1948) and the data processing inequality (DPI) (Pippenger, 1988). Shannon’s theory fails to account for scenarios where $X$ contains less usable information than the mutual information $I(X;Y)$ due to encryption. Similarly, DPI cannot explain how model family $\mathcal{V}$ acquires additional information through computational constraints or advanced representation learning. Two key factors demonstrate this limitation; (1) computational constraints prevent input data from fully representing ideal world knowledge (Xu et al., 2020); (2) advanced language models can extract meaningful features from incomplete representations, achieving progressive information gains during computation (LeCun et al., 2015; Goldfeld and Greenewald, 2021).
Recent work has introduced two frameworks adopting $\mathcal{V}$-usable information to capture these phenomena. The first framework to capture the $\mathcal{V}$-usable information is called predictive $\mathcal{V}$-information (Xu et al., 2020). The predictive $\mathcal{V}$-information measures how much information can be extracted from X about Y when constrained to model family $\mathcal{V}$, written as $I_{\mathcal{V}}(X\rightarrow Y)$. The greater the $I_{\mathcal{V}}(X\rightarrow Y)$, the easier the dataset is for $\mathcal{V}$. While predictive $\mathcal{V}$-information provides an aggregate measure of informativeness of computational functions or dataset difficulty, pointwise $\mathcal{V}$-information (Ethayarajh et al., 2022) measures usable information in individual instances with respect to a given dataset distribution, written as $\text{PVI}(x\rightarrow y)$. The higher the PVI, the easier the instance is for $\mathcal{V}$, under the given distribution.
We first define predictive conditional $\mathcal{V}$-entropy to introduce the predictive $\mathcal{V}$-information. We follow the formal notations, defined in Xu et al. (2020):
Definition 2.1
(Xu et al., 2020) Let predictive family $\mathcal{V}\subseteq\Omega=\{f:\mathcal{X}\cup\varnothing\rightarrow P(\mathcal{Y})\}$, where $X$ and $Y$ are random variables with sample space $\mathcal{X}$ and $\mathcal{Y}$, and $P(\mathcal{Y})$ be the set of all probability measures on $\mathcal{Y}$ over the Borel algebra on $\mathcal{X}$. The predictive conditional $\mathcal{V}$-entropy is defined as
$H_{\mathcal{V}}(Y|X)=\text{inf}_{f\in\mathcal{V}}\mathbb{E}_{x,y\sim X,Y}[-\log_{2}f[X](Y)].$ (1)
The conditional $\mathcal{V}$-entropy is given a random variable $X$ as side information, so the function $f[X](Y)$ produces probability distributions over the output $Y$ based on the side information $X$. Suppose $\varnothing$ denote a null input that provides no information about $Y$. Note that $\varnothing\notin X$. The predictive family $\mathcal{V}$ is a subset of all possible mappings from $X$ to $P(Y)$ that satisfies optional ignorance; whenever the $P$ predicts the outcome of $Y$, it has the option to ignore side information, $X$. That is,
$H_{\mathcal{V}}(Y|\varnothing)=\text{inf}_{f\in\mathcal{V}}\mathbb{E}_{y\sim Y}[-\log_{2}f[\varnothing](Y)].$
This is identical to the classic $\mathcal{V}$-entropy, denoted as $H_{\mathcal{V}}(Y)$. We additionally specify the notation $\varnothing$ because the conditional $\mathcal{V}$-information given null input is crucial to measure how the existence of $X$ affect $\mathcal{V}$ to obtain the relevant information. The entropy estimation specifies the infinite functions $f$ are in $\mathcal{V}$ as Xu et al. (2020) illustrate that the predictive family in theory does not take into account the computational constraints.
Definition 2.2
(Xu et al., 2020) Let $X$, $Y$ denote random variables with sample space $\mathcal{X}\times\mathcal{Y}$, and $\mathcal{V}$ be a predictive model or function family. Then the predictive $\mathcal{V}$-information from $X$ to $Y$ is defined as
$I_{\mathcal{V}}(X\rightarrow Y)=H_{\mathcal{V}}(Y|\varnothing)-H_{\mathcal{V}}(Y|X).$ (2)
Definition 2.3
(Ethayarajh et al., 2022) Given random variables $X,Y$ and a predictive family $\mathcal{V}$, the pointwise $\mathcal{V}$-information (PVI) of an instance $(x,y)$ is
$I_{\mathcal{V}}(x\rightarrow y)=-\log_{2}{f^{\prime}[\varnothing](y)}+\log_{2}{f^{\prime}[x](y)}.$ (3)
Ethayarajh et al. (2022) have extended the Equation 2 to estimate the difficulty of point-wise instances for the predictive family $\mathcal{V}$. Most neural networks fine-tuned to fit label information $Y$ meet the definition of the predictive family $\mathcal{V}$ here. If $\mathcal{V}$ were, for instance, the BERT function family, $f^{\prime}[X]$ and $f^{\prime}[\varnothing]$ would be the models after finetuning BERT with and without the input $X$ respectively. Higher PVI means that the instance is easy for $\mathcal{V}$ while lower PVI means difficult among the given distribution. This comes from the intuition that predicting minority instances expects $\mathcal{V}$ to require more side information of $X$ to understand the instances.
3 Layer-Wise Usable Information
We use the $\mathcal{V}$-usable information framework to estimate the predictive uncertainty in natural language generation. Given a set of context $C$, language models generate answers in a free-form text to questions $Q$. In this framework, we consider a formally defined predictive family $\mathcal{V}$ as one of pre-trained generative language models that are consist of a number of layers $\mathcal{L}$. Each layer $\ell$ in language models creates different computations. Under the hypothesis that each layer $\ell$ has a different amount of executable information, we propose layer-wise usable information ($\mathcal{L}$I).
Let predictive family $\mathcal{L}\subseteq\Omega=\{f^{(\ell)}:\mathcal{C}\cup\varnothing\rightarrow P(\mathcal{Q})\}$, where $C$ and $Q$ are random variables with sample spaces $\mathcal{C}$ and $\mathcal{Q}$, and $P(\mathcal{Q})$ be the set of all probability measures on $\mathcal{Q}$ over the Borel algebra on $\mathcal{C}$. The function $f^{(\ell)}$ indicates a function of $\ell$, one of layers within the predictive family $\mathcal{L}$ (or $\mathcal{V}$), i.e., a pre-trained language model. The $\mathrm{range}(f)$ is a vocabulary size of $\mathcal{L}$. Unlike prior work, a function $f$ is not fine-tuned on training data.
Definition 3.1.
(Predictive conditional $\ell$-entropy). Let $q$ denotes a question token in $Q$. The predictive conditional $\ell$-entropy is
$H_{\ell}(Q|C)=\mathbb{E}_{q\sim Q}[-\log_{2}p(q_{t}|q_{<t},C)].$ (4)
$H_{\ell}(Q|\varnothing)=\mathbb{E}_{q\sim Q}[-\log_{2}p(q_{t}|q_{<t},\varnothing)].$
Given a layer $\ell$ in a generative language model, we define $f:{C\cup\varnothing}\rightarrow P(\mathcal{Q})$ as a mapping function from context tokens (or their absence) to probability distributions over question token sequences. These formulations characterize the model’s capacity to generate probabilistic distributions over question tokens, with the fundamental distinction lying in the conditioning on context tokens $C$ or their absence $\varnothing$. This mathematical framework elucidates the contextual influence on token-level probability assignments within the generative process.
Definition 3.2.
(Predictive $\mathcal{L}$-information). Then the predictive $\mathcal{L}$-information from $C$ to $Q$ is defined as
$\mathcal{L}I(c\rightarrow q)$ (5)
$=\sum_{\ell\in\mathcal{L}}I_{\ell}(c\rightarrow q)$
$=\sum_{\ell\in\mathcal{L}}H_{\ell}(Q|C)-H_{\ell}(Q|\varnothing)$
=$=\sum_{\ell\in\mathcal{L}}\{-\log_{2}p(q_{t}|q_{<t},\varnothing)+\log_{2}p(q_{t}|q_{<t},c)\}.$
To measure the information of $C$ to $P(Q)$, we compute the difference between the entropies with contextual information $C$ and without contextual information $\varnothing$ across layers $\mathcal{L}$. Algorithm 1 provides detailed steps for the calculation.
The useful properties of $\mathcal{L}I$ follows that of $\mathcal{V}$ information:
  • Non-negativity: $\mathcal{LI}(C\rightarrow Q)\geq 0$
  • Independence: If $C$ is not relevant to $Q$, $\mathcal{LI}(C\rightarrow Q)=0$.
  • Monotonicity: If predictive families $\mathcal{U}\subseteq\mathcal{L}$, then $\mathcal{LI_{U}}(Q)\geq\mathcal{LI_{L}}(Q)$ and thus $\mathcal{LI_{U}}(Q|C)\geq\mathcal{LI_{L}}(Q|C)$.
Algorithm 0 $\mathcal{L}$VI
Input: data $\mathcal{D}=\{(\text{context}\ c_{i},\text{question}\ q_{i})\}_{i=1}^{m}$, model $\mathcal{L}$.
[…]
  $f^{\ell}[c]\leftarrow$ Inference with $\mathcal{L}$ on $\mathcal{D}$
  $f^{\ell}[\varnothing]\leftarrow$ Inference with $\mathcal{L}$ on $\{(\varnothing,q_{i})\ |\ (c_{i},q_{i})\in\mathcal{D}\}$
  $H_{\mathcal{L}}(Q|\varnothing),H_{\mathcal{L}}(Q|C)\leftarrow 0,0$
  for $(l)\in\mathcal{L}\text{ layers }\text{in model}\mathcal{L}$ do
     $H_{\ell}(Q|\varnothing)\leftarrow H_{\ell}(Q|\varnothing)-\mathbb{E}_{q\sim Q}[\log_{2}p(q_{t}|q_{<t},\varnothing)]$
     $H_{\ell}(Q|C)\leftarrow H_{\ell}(Q|C)-\mathbb{E}_{q\sim Q}[\log_{2}p(q_{t}|q_{<t},C)]$
     $\mathcal{L}\text{I}(c_{i}\to q_{i})\leftarrow H_{\ell}(Q|\varnothing)-H_{\ell}(Q|C)$
  end for
  $\mathcal{L}I(C\to Q)=\frac{1}{m}\sum_{i}\mathcal{L}\text{I}(c_{i}\to q_{i})\newline \text{\hskip 80.00012pt}=H_{\mathcal{L}}(Q)-H_{\mathcal{L}}(Q|C)$
Algorithm 1 $\mathcal{L}$I can be calculated with two conditional inferencing strategies (without fine-tuning).
3.1 Assumptions
$\mathcal{L}$I measures how confident the language models are in their generated responses. To measure the $\mathcal{L}$I, we assume the pre-trained language models are optimal distributions for two random variables $C$ and $Q$, contextual information and question prompts, respectively. We also assume that the difference between the true $\mathcal{V}$-information, derived from the optimal distribution, and the empirical $\mathcal{V}$-information, obtained through in-context prompting procedures with pre-trained language models, is negligible.
3.2 Data Processing Inequality In Language Models
The Data Processing Inequality (DPI) states that $I(X;Y)\geq I(X;Z)$, given a Markov chain, $X\rightarrow Y\rightarrow Z$. Because conditional distribution of $Z$ depends only on $Y$ and is conditionally independent from $X$, the additional processing of $Y$ decreases the information about $X$ for $Z$ (Pippenger, 1988; Feder, 1989; Beaudry and Renner, 2011). $I(X;Z)=I(X;Y)$ if and only if $Z$ and $Y$ have the same information about $X$. However language models violate the statement. Language models process the input tokens through layers to extract features that are increasingly useful for inferring desirable responses (LeCun et al., 2015; Goldfeld and Greenewald, 2021). More formally, denoting $\gamma$ as the feature extraction algorithm and $\mathcal{V}$ as a class of predictive language models, we can state that $I_{\mathcal{V}}(\gamma(X)\rightarrow Y)>I_{\mathcal{V}}(X\rightarrow Y)$. Yet we empirically argue that language models do not always increase or decrease the information quantity over feature extraction but fluctuate it within the mechanism of updating multiple layers, as illustrated in Figure 2(a). $\mathcal{L}$I covers the inconsistency of both obeying and violating the DPI in layer-wise functions of generative language models by aggregating the layer-by-layer usable information quantities.
3.3 Implications
Using layer-wise usable information ($\mathcal{L}$I), we contribute to the following accomplishments:
  • Detection for unanswerable questions by computing $\mathcal{L}I(c\rightarrow q)$ in datasets $c\in C,q\in Q$} for the same $\mathcal{L}$: we classify questions that lack sufficient usable information as unanswerable, likely to be inaccurate responses (Fig. 3(b)).
  • Evaluation on different prompts with $Q$ for $\mathcal{L}$ by estimating $\mathcal{L}I(C\rightarrow Q^{\prime})$ for the same $\mathcal{L}$ and $C$ but different $Q^{\prime}$ (Figs. 1 4 and 5).
  • Analysis of model performance estimating $\mathcal{L}I(C\rightarrow Q)$: we can comprehensively predict the overall performance of the language model across different scenarios and contexts (Figs. 1 and 6).
4 Related Work
Existing work has analyzed the model’s capability to detect unanswerable questions from three main perspectives. One is self-evaluation which allows language models to generate probabilistic scores of how much the models believe their answers are trustworthy (Kadavath et al., 2022a; Yin et al., 2023) as we get access to advanced, well-calibrated models that can generate reliable results. The second perspective involves identifying the subspace of the model that is specifically responsible for answerability (Slobodkin et al., 2023). The third approach uses label information to train LLMs on whether questions are answerable, employing methods such as instruction-tuning or calibration (Jiang et al., 2021; Kapoor et al., 2024). While these studies suggest that LLMs can learn to express their confidence in responses when provided with additional information, they rely heavily on external calibration or fine-tuning. The outcome depends on the quality of the additional information or that of the annotation work. Unlike prior work, our investigation does not require label annotations to fine-tune classifiers or calibration tools to detect model confidence in their generated answers or ambiguous, unanswerable questions. We aim to obtain a computationally feasible method that applies to universal large language models.
5 Experiments
We demonstrate that layer-wise usable information ($\mathcal{L}$I) is an effective way to detect unanswerable questions or capture the ambiguity of prompts for large language models.
5.1 Experimental Setup
Evaluation Metric. We classify unanswerable questions based on uncertainty scores. We evaluate uncertainty under the assumption that we assess whether to trust a model’s generated response in a given context, i.e., deciding whether to accept an answer to a question. Our primary metric for this assessment is the area under the receiver operating characteristic curve (AUROC). The AUROC metric measures how well the model can differentiate between correct and incorrect answers, or between answerable and unanswerable questions in our case. Higher AUROC scores indicate better performance, with a perfect score of 1 representing optimal uncertainty estimation, while a score of 0.5 represents random uncertainty.
We choose to use the AUROC as it suits well for evaluating uncertainty in free-form text responses, as opposed to calibration measures like the Brier score, frequently used in classification tasks or multiple-choice question answering. The Brier score requires calculating the total probability mass assigned to all possible tokens of a correct answer sequences. This causes the task to be intractable in free-form text settings where probabilities with respect to meanings are unavailable. Therefore we use the AUROC to capture the uncertainty associated with the model’s outputs more accurately, and classify the unanswerable questions. One exception is model answers that we simply match the lexical words by instructing models to generate "Yes" or "No" to additional question prompts asking if they can answer the question based on the context.
Baselines. We evaluate our method against several benchmarks, including model-generated answers, P(True) (Kadavath et al., 2022b), predictive token entropy and normalized entropy (Malinin and Gales, 2020), semantic entropy (Farquhar et al., 2024), and pointwise $\mathcal{V}$-information (PVI) (Ethayarajh et al., 2022) on the first and the last layers respectively. Model-generated answers are raw responses by models. P(True) measures the probability that a LLMs predict the next token as ‘True’ when provided with few-shot prompts that compare a primary answer to various alternative answers. Predictive entropy is calculated by conditional entropy over the output distribution. The predictive normalized entropy divides the joint joint log-probability of each sequence by their length. We use a single model to meausre the normalized entropy, following the setups by (Kuhn et al., 2023b). Semantic entropy follows the confabulation mechanism to classify unanswerable questions. PVI is to measure difficult datapoints. We assume that difficult instances for language models are likely to be unanswerable questions.
Models. We use Llama3 (Dubey et al., 2024) and Phi3 models (Abdin et al., 2024). We vary the size of the models between 3.8B, 8B, and 14B parameters. We report our headline results using the most computationally efficient model, with 3.8B parameters unless we notify otherwise. In all cases we use only a single unmodified model since recent foundation models are not practical to modify the architectures and are often too costly to fine-tune them on datasets. Above all, we are interested in investigating internal language model behaviors than simply achieving optimal performance results. Hence we use them in their pre-trained form.
Datasets. We use Conversational Question Answering Challenge dataset (CoQA) (Reddy et al., 2019) and Question Answering In Context (QuAC) (Choi et al., 2018) as question-answering tasks, where the model responds to questions using information from a supporting context paragraph. Our experiments are conducted on the development set, which contains approximately 8,000 questions. We also use CondaQA (Ravichander et al., 2022) which features 14,182 question-answer pairs with over 200 unique negation cues in addition to CoQA and QuAC to evaluate how trustworthy the $\mathcal{L}$I is to detect unanswerable questions. Given that goal, we employ a 1-to-1 ratio of answerable to unanswerable questions for a clear performance evaluation.
5.2 Do $\mathcal{L}$I scores capture unanswerable questions?
(a) Performance comparison of $\mathcal{L}$I scores with other baselines.
(b) Performance comparison on different types of models.
Figure 3: Performance of (un)answerability detection.
Across all three benchmark datasets — CoQA, QuAC, and CondaQA — $\mathcal{L}$I estimation consistently outperforms all baseline methods in classifying answerable and unanswerable questions as we show in Figure 3(a). We observe in Figure 3(b) that $\mathcal{L}$I is consistently good at detecting unanswerable questions over different model settings. Considering the unique purpose of detecting unanswerable questions, prior uncertainty estimation may not be proper approaches. For example, semantic entropy (SE) shows solid performances in diverse tasks but (un)answerability tasks, as the models are restricted to generating little variations in answers.
To further investigate the correlation between low $\mathcal{L}$I scores and unanswerable questions, we use rejection rates and evaluate the performances. The rejected answers are counted towards unanswerable questions for this evaluation. We observe that the AUROC of $\mathcal{L}$I goes up by rejecting data points with low percentile of $\mathcal{L}$I scores. That is, the low $\mathcal{L}$I scores are indicative of unanswerable questions, contributing positively to the AUROC of the $\mathcal{L}$I metric.
We observe that Llama-3 model (8B) performs better than Phi-3 medium (14B) because Llama models are more cautious to generate answers than Phi-3 models. Llama shows higher ratio of generating ’No’ to ’Yes’ for the questions asking if they can answer the question based on the given context than both Phi-3 mini (3.8B) and Phi-3 (14B) do. Because we present the performances on dataset with the same ratio of answerable to unanswerable questions, Llama-3 (8B) shows high AUROC scores. Despite the biased preferences in choosing answers by LLMs, $\mathcal{L}$I scores consistently outperform the other methods, being less sensitive to those changes in settings.
We observe Figure 4 that $\mathcal{L}$I shows apparent gap in answerable and unanswerable questions regardless of the existence of prompts, while the prompts obviously increase the overall $\mathcal{L}$I scores, leading LLMs to be more certain about their decision making.
5.3 Do $\mathcal{L}$I scores indicate the ambiguity of prompts?
$\mathcal{L}$I scores capture the ambiguity of prompts. We previously observed Figure 1 that both model performances and $\mathcal{L}$I scores show apparent gaps between those with and without instruction prompts. While we observe that the $\mathcal{L}$I scores show strong correlations with the model performances, Figure 2 supported our claims that instruction prompts increase overall $\mathcal{L}$I scores on each dataset. We also observed Figure 4 that $\mathcal{L}$I scores capture advantages of prompts across all the benchmark datasets while consistently producing lower scores on unanswerable questions than on answerable questions.
In Figure 5, we observe that $\mathcal{L}$I scores show comprehensive understanding of different levels of ambiguity stemming from the given instruction prompts. We observe the consistent yet more refined phenomena with answerable and unanswerable questions while their mean scores show that LLMs are most uncertain when the instruction prompts are never given, slightly certain given open-ended prompts with an option for responding that the answer is unknown in the context, and most certain when the prompts instruct LLMs to answer either yes or no. Both by intuition and model performances, we verify that $\mathcal{L}$I scores are valid showing the different model certainty depending on the instruction prompts settings.